# Double Fairness Policy Learning: Integrating Action Fairness and Outcome Fairness in Decision-making

### v1

@@ 5 Numerical Studies
Data generating procedure. We consider the following data-generating process, which is similar to that in Section F.2.1 of Wang et al. (2025). The sensitive variable is given by $S_{i}\sim\mbox{Bernoulli}(0.35)$, and we consider the case in which the covariates are two-dimensional and are generated by $X_{1i}=0.45S_{i}+\varepsilon_{1i}$, and $X_{2i}=0.85S_{i}+\varepsilon_{2i}$, where $\{\varepsilon_{1i}\}_{i}$ and $\{\varepsilon_{2i}\}_{i}$ are independent error terms drawn from a standard Gaussian distribution. The action is sampled using $A_{i}\sim\mbox{Bernoulli}(\operatorname*{expit}(-0.5+S_{i}+X_{1i}-X_{2i}))$, where $\operatorname*{expit}(t)=\exp(t)/(1+\exp(t))$. Finally, we consider the case where $R=R^{(1)}=R^{(2)}$, such that $R$ is one-dimensional. In this setting, the reward is generated according to: $R_{i}=1.5S_{i}+0.9X_{1i}+0.8X_{2i}+A_{i}(1-0.8S_{i}-0.5X_{1i}-0.5X_{2i})+\epsilon_{i},$ where $\{\epsilon_{i}\}_{i}$ are independent error terms drawn from a standard Gaussian distribution. Under this setting, the global optimal decision rule is a linear decision rule $\mathds{1}(1-0.8S_{i}-0.5X_{1i}-0.5X_{2i}>0)$.
Competing methods. We compare our proposed method with the standard policy learning approach, and the approach for finding Pareto policies in Viviano and Bradic (2023). Specifically, the standard policy learning approach estimates the optimal policy as
@@
As for the approach outlined in Viviano and Bradic (2023), it focuses solely on single fairness policy learning tasks, either outcome fairness or action fairness. To build on this, we first consider the original method based on the action fairness metric (VB1) and the outcome fairness metric (VB2). Moreover, we consider an adaptive version (ADVB), which has been tailored to accommodate the double fairness setting. Specifically, the ADVB method is adapted to our proposed double fairness setting utilizing the linear scalarization method to derive the resulting policy. As in our proposed procedure, it involves two steps: first, the estimated Pareto set $\widehat{\Pi}_{ADVB}$ is obtained using linear scalarization, i.e.,
@@ 6.1 Motor Insurance Liability Claims Data
In this section, we illustrate our proposed method to a Belgian motor third-party liability dataset from the CASdatasets (Dutang and Charpentier, 2019). Similar to Xin and Huang (2024), we use gender as the sensitive variable. In the offline dataset, we classify customers as low-risk or high-risk based on the bonus-malus (BM) scale, following common practice in the insurance literature, where drivers are often categorized as good or bad based on their BM score (Frangos and Vrontos, 2001). The BM scale ranges from 0 to 22, with higher values indicating greater risk. We define customers with BM greater than five as high-risk and the rest as low-risk. The dataset consists of 18,276 observations, with 28% being female and 35% classified as high-risk. The non-sensitive covariates include the number of claims, policyholder age, coverage plan, if the vehicle is part of a fleet, and the coverage duration. Since the dataset does not contain premium information, we generate the pure premium amount using the frequency-severity approach, as outlined by Baumann and Loi (2023). This approach sets pure premiums based on the predicted risk. To better reflect real-world pricing, we further adjust the final premium based on the coverage plan and gender. See Section 8.4 in Supplemental Materials for more details. Accordingly, the benefit or welfare received by the customer is the negative of the reward. Hence, it is essential to consider both action and outcome fairness in this setting.
@@ 6.2 Entrepreneurship Training Data
In this subsection, we illustrate our proposed method to the entrepreneurship training data for North America undergraduate students from Lyons and Zhang (2017). This study examined whether entrepreneurship programs have varying impacts on certain minority groups, such as females or non-Caucasians. Focusing solely on the outcome fairness criterion, they found that the program’s short-term effect on minorities is minimal. The goal in our study is to identify the policy that maximizes the expected subsequent entrepreneurial activity for the population while keeping the action and value unfairness as low as possible.
@@ 8.1 Limitation of the Linear Scalarization Approach
(a) Linear scalarization.
(b) Weighted Tchebycheff scalarization.
Figure 7: Linear scalarization approach cannot identify Pareto polices in the non-convex regions of the Pareto front, as these solutions lack supporting hyperplanes (Miettinen, 1999). In contrast, weighted Tchebycheff scalarization can find all Pareto policies, regardless of the shape of the objective space. Figure adapted from Lin et al. (2024).
@@ 8.4 Further Details on Premium in the Insurance Data
In the offline dataset, we classify customers as low-risk or high-risk based on the BM scale, following common practice in the insurance literature, where drivers are often categorized as good or bad based on their BM score (Frangos and Vrontos, 2001). The BM scale ranges from 0 to 22, with higher values indicating greater risk. We define customers with BM greater than five as high-risk and the rest as low-risk. Since the dataset does not contain premium information, we generate the pure premium amount using the frequency-severity approach, as outlined by Baumann and Loi (2023). This approach sets pure premiums based on the predicted risk. To better reflect real-world pricing, we further adjust the final premium based on the coverage plan and gender. Specifically, the premium is generated according to
@@ 9.3 Proof of Proposition 3
To show all of the elements in $\Pi_{p}$ is Pareto optimal solution, see Theorem 5.4.1 in Miettinen (1999); while the other direction can be proved using Theorem 5.4.2 in Miettinen (1999). ∎
@@ 9.6 Proof of Theorem 3: Equal Opportunity
by Assumption 4.4. Using standard empirical process techniques (see, e.g., Van Der Vaart and Wellner (1996)), the second term is of order
@@
are still VC-type, because squaring is a monotone Lipschitz map on $[0,1]$ and products of bounded VC-type classes remain VC-type (see Lemma 2.6.20 in Van Der Vaart and Wellner (1996)). Thus, a direct application of Dudley’s entropy integral bound yields that
@@
i.e., the oracle estimator of $\Delta_{2}(\pi)$, and let $\Delta_{2i}=(f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i}))^{2}$. Applying the symmetrization technique (Van Der Vaart and Wellner, 1996), we obtain

### v2

@@ 5 Numerical Studies
Data generating procedure. We consider the following data-generating process, which is similar to that in Section F.2.1 of wang2025counterfactually. The sensitive variable is given by $S_{i}\sim\mbox{Bernoulli}(0.35)$, and we consider the case in which the covariates are two-dimensional and are generated by $X_{1i}=0.45S_{i}+\varepsilon_{1i}$, and $X_{2i}=0.85S_{i}+\varepsilon_{2i}$, where $\{\varepsilon_{1i}\}_{i}$ and $\{\varepsilon_{2i}\}_{i}$ are independent error terms drawn from a standard Gaussian distribution. The action is sampled using $A_{i}\sim\mbox{Bernoulli}(\operatorname*{expit}(-0.5+S_{i}+X_{1i}-X_{2i}))$, where $\operatorname*{expit}(t)=\exp(t)/(1+\exp(t))$. Finally, we consider the case where $R=R^{(1)}=R^{(2)}$, such that $R$ is one-dimensional. In this setting, the reward is generated according to: $R_{i}=1.5S_{i}+0.9X_{1i}+0.8X_{2i}+A_{i}(1-0.8S_{i}-0.5X_{1i}-0.5X_{2i})+\epsilon_{i},$ where $\{\epsilon_{i}\}_{i}$ are independent error terms drawn from a standard Gaussian distribution. Under this setting, the global optimal decision rule is a linear decision rule $\mathds{1}(1-0.8S_{i}-0.5X_{1i}-0.5X_{2i}>0)$.
Competing methods. We compare our proposed method with the standard policy learning approach, and the approach for finding Pareto policies in viviano2023fair. Specifically, the standard policy learning approach estimates the optimal policy as
@@
As for the approach outlined in viviano2023fair, it focuses solely on single fairness policy learning tasks, either outcome fairness or action fairness. To build on this, we first consider the original method based on the action fairness metric (VB1) and the outcome fairness metric (VB2). Moreover, we consider an adaptive version (ADVB), which has been tailored to accommodate the double fairness setting. Specifically, the ADVB method is adapted to our proposed double fairness setting utilizing the linear scalarization method to derive the resulting policy. As in our proposed procedure, it involves two steps: first, the estimated Pareto set $\widehat{\Pi}_{ADVB}$ is obtained using linear scalarization, i.e.,
@@ 6.1 Motor Insurance Liability Claims Data
In this section, we illustrate our proposed method to a Belgian motor third-party liability dataset from the CASdatasets (dutang2019casdatasets). Similar to xin2024antidiscrimination, we use gender as the sensitive variable. In the offline dataset, we classify customers as low-risk or high-risk based on the bonus-malus (BM) scale, following common practice in the insurance literature, where drivers are often categorized as good or bad based on their BM score (frangos2001design). The BM scale ranges from 0 to 22, with higher values indicating greater risk. We define customers with BM greater than five as high-risk and the rest as low-risk. The dataset consists of 18,276 observations, with 28% being female and 35% classified as high-risk. The non-sensitive covariates include the number of claims, policyholder age, coverage plan, if the vehicle is part of a fleet, and the coverage duration. Since the dataset does not contain premium information, we generate the pure premium amount using the frequency-severity approach, as outlined by baumann2023fairness. This approach sets pure premiums based on the predicted risk. To better reflect real-world pricing, we further adjust the final premium based on the coverage plan and gender. See Section 8.4 in Supplemental Materials for more details. Accordingly, the benefit or welfare received by the customer is the negative of the reward. Hence, it is essential to consider both action and outcome fairness in this setting.
@@ 6.2 Entrepreneurship Training Data
In this subsection, we illustrate our proposed method to the entrepreneurship training data for North America undergraduate students from lyons2017impact. This study examined whether entrepreneurship programs have varying impacts on certain minority groups, such as females or non-Caucasians. Focusing solely on the outcome fairness criterion, they found that the program’s short-term effect on minorities is minimal. The goal in our study is to identify the policy that maximizes the expected subsequent entrepreneurial activity for the population while keeping the action and value unfairness as low as possible.
@@ 8.1 Limitation of the Linear Scalarization Approach
((a)) Linear scalarization.
((b)) Weighted Tchebycheff scalarization.
Figure 7: Linear scalarization approach cannot identify Pareto polices in the non-convex regions of the Pareto front, as these solutions lack supporting hyperplanes (miettinen1999nonlinear). In contrast, weighted Tchebycheff scalarization can find all Pareto policies, regardless of the shape of the objective space. Figure adapted from lin2024smooth.
@@ 8.4 Further Details on Premium in the Insurance Data
In the offline dataset, we classify customers as low-risk or high-risk based on the BM scale, following common practice in the insurance literature, where drivers are often categorized as good or bad based on their BM score (frangos2001design). The BM scale ranges from 0 to 22, with higher values indicating greater risk. We define customers with BM greater than five as high-risk and the rest as low-risk. Since the dataset does not contain premium information, we generate the pure premium amount using the frequency-severity approach, as outlined by baumann2023fairness. This approach sets pure premiums based on the predicted risk. To better reflect real-world pricing, we further adjust the final premium based on the coverage plan and gender. Specifically, the premium is generated according to
@@ 9.3 Proof of Proposition 3
To show all of the elements in $\Pi_{p}$ is Pareto optimal solution, see Theorem 5.4.1 in miettinen1999nonlinear; while the other direction can be proved using Theorem 5.4.2 in miettinen1999nonlinear. ∎
@@ 9.6 Proof of Theorem 3: Equal Opportunity
by Assumption 4.4. Using standard empirical process techniques (see, e.g., van1996weak), the second term is of order
@@
are still VC-type, because squaring is a monotone Lipschitz map on $[0,1]$ and products of bounded VC-type classes remain VC-type (see Lemma 2.6.20 in van1996weak). Thus, a direct application of Dudley’s entropy integral bound yields that
@@
i.e., the oracle estimator of $\Delta_{2}(\pi)$, and let $\Delta_{2i}=(f^{\pi}(1,X_{i})-f^{\pi}(0,X_{i}))^{2}$. Applying the symmetrization technique (van1996weak), we obtain